OpenStudy (anonymous):

You go back-to-school shopping and get 3 pairs of shoes, 7 shirts, and 4 pairs of pants for $410. Your friend got 4 pairs of shoes and 6 pairs of pants for $380. Your other friend got 5 shirts and 3 pairs of pants for $190. Using x for shoes, y for shirts, and z for pants; write a system of equation to solve for the price of each item.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

3x+7y+4z= 410 4x+6z= 380 5y+3z=190

OpenStudy (anonymous):

Is it correct?

jimthompson5910 (jim_thompson5910):

let me check

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

Everything looks great. Nice work

OpenStudy (anonymous):

Thanks for helping me! I'm starting to get understand it more and more.

jimthompson5910 (jim_thompson5910):

I'm glad you are